OpenStudy (anonymous):

A body moves along a horizontal line according to s=f(t)=t^3-9t^2+24t a) when is s increasing and when it is decreasing? b) when v is increasing and when it is decreasing?

OpenStudy (richyw):

\[\frac{ds}{dt} = 3t^2-18t+24\]\[\frac{d^2s}{dt^2} = 6t-18\] a) when \(\frac{ds}{dt}>0\), s is increasing, when \(\frac{ds}{dt}<0\), s is decreasing b) when \(\frac{d^2s}{dt^2}>0\), v is increasing, when \(\frac{d^2s}{dt^2}>0\), v is decreasing

OpenStudy (richyw):

oops the last statement in b should have a < sign (just like in a)